OpenStudy (anonymous):

If tanx= -5/12 and x is in the 2nd quadrant, find sin2x. A) 10/13 B) 120/169 C) 0 D) -120/169

OpenStudy (anonymous):

sin 2x=2sinx cos x in the second quadrant sine is positive and cosine is negative so sin2x is negative.

OpenStudy (anonymous):



OpenStudy (anonymous):

from the figure sin x=5/13 and cos x=-12/13

OpenStudy (anonymous):

this sin 2x=2*(5/13)*(-12/13)=-120/169

OpenStudy (anonymous):

option D is correct

OpenStudy (anonymous):

@aminaj_xoxo